# A Closer Look on Memorization in Tabular Diffusion Model: A Data-Centric Perspective

### v1

@@ 6.3 DynamicCut with TabCutMix (DynamicCutMix): Combined Effect and Evaluation
cell21 = r=6,
cell81 = r=6,
cell141 = r=6,
cell201 = r=6,
cell261 = r=6,
hline1-2,8,14,20,26 = -,
& Tagging Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Default
- CTGAN 0.63 - 68.68 0.21 68.95 1.70 16.49 0.55 85.04 0.93 77.05 2.47 61.89 4.89 49.45 0.80
CTGAN CTGAN+DCM 0.07 72.37 0.39 71.72 0.74 16.64 0.20 85.78 0.70 76.85 1.62 57.98 3.09 48.48 0.24
TabDDPM CTGAN+DCM 0.32 71.41 0.89 68.93 0.94 16.09 0.59 84.42 1.03 75.38 1.54 56.85 4.31 50.49 1.12
- TVAE 0.43 - 72.24 0.36 82.97 0.37 20.57 0.42 89.37 0.54 83.36 0.99 52.63 0.37 51.33 0.96
TVAE TVAE+DCM 0.21 76.27 0.25 85.49 0.41 21.66 0.19 90.36 0.25 86.76 5.13 54.81 0.79 51.07 1.35
TabDDPM TVAE+DCM 0.27 73.82 0.40 80.89 1.12 21.47 1.28 89.58 0.19 78.24 1.28 53.27 4.92 50.92 2.49
Shoppers
- CTGAN 1.15 - 83.22 1.30 85.26 5.17 26.67 1.66 77.73 0.50 86.48 0.69 67.18 5.84 48.54 0.74
CTGAN CTGAN+DCM 0.56 85.51 0.46 83.51 1.60 26.19 0.82 78.23 0.30 86.36 0.47 68.85 2.45 50.64 1.11
TabDDPM CTGAN+DCM 0.43 83.52 0.53 80.21 2.75 25.70 1.22 77.49 1.24 86.40 0.62 66.03 7.67 50.47 2.27
- TVAE 0.77 - 86.65 0.48 50.32 2.98 11.16 1.33 74.75 0.92 77.56 1.11 21.29 3.28 43.31 0.65
TVAE TVAE+DCM 1.33 88.32 0.29 51.12 2.46 11.98 0.58 75.02 0.97 79.02 0.12 22.99 2.12 49.98 2.55
TabDDPM TVAE+DCM 0.95 88.43 0.17 50.43 3.45 12.08 2.01 74.82 0.25 78.37 1.18 22.57 3.68 47.26 1.81
Adult
- CTGAN 0.62 - 88.64 0.32 78.15 3.66 26.27 0.67 82.43 0.90 82.83 0.93 63.64 2.74 49.14 0.27
CTGAN CTGAN+DCM 0.88 88.45 0.29 79.31 1.04 23.31 1.42 82.23 0.23 82.80 2.27 63.46 3.43 48.13 0.33
TabDDPM CTGAN+DCM 0.32 87.87 0.37 72.40 4.17 25.56 1.45 81.42 1.91 81.15 1.23 58.84 2.46 50.00 0.19
- TVAE 0.41 - 88.61 0.49 92.34 2.19 29.90 1.00 82.71 0.45 79.06 0.90 48.67 2.70 48.76 0.25
TVAE TVAE+DCM 0.45 88.12 0.32 87.54 1.24 31.54 0.38 83.45 0.84 78.57 0.68 54.85 5.05 48.81 2.10
TabDDPM TVAE+DCM 0.21 89.23 0.38 85.50 4.09 30.23 3.81 82.36 1.96 78.80 2.50 54.48 8.01 49.57 0.35
@@ A.1.1 Overall Performance Comparison
cell21 = r=4,
cell61 = r=4,
cell101 = r=4,
cell141 = r=4,
cell181 = r=4,
hline1-2,6,10,14,18 = -,
& Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Shoppers
TabDDPM 0.31 - 92.17 0.32 93.16 1.58 52.57 1.30 97.08 0.46 92.92 3.27 86.74 0.63 51.36 0.63
TabDDPM+DCM 5% 1.76 93.30 0.22 94.49 1.87 49.07 2.35 95.98 1.14 92.04 1.81 84.66 3.35 51.28 1.28
TabDDPM+DCM 10% 2.83 91.46 0.74 93.25 3.87 46.15 4.42 92.32 5.00 89.11 4.95 80.56 4.81 49.74 3.34
TabDDPM+DCM 20% 0.40 91.58 0.51 91.66 2.77 52.98 1.99 95.30 1.88 91.83 1.35 82.69 4.29 52.44 0.33
Adult
TabDDPM 0.18 - 91.09 0.07 93.58 1.99 51.52 2.29 98.84 0.03 97.78 0.07 94.63 1.19 51.56 0.34
TabDDPM+DCM 5% 1.41 90.91 0.26 95.73 2.92 45.56 2.63 98.08 0.48 96.14 1.50 90.21 3.35 51.06 0.56
TabDDPM+DCM 10% 0.57 90.81 0.16 95.23 1.14 45.76 1.02 98.27 0.32 96.71 0.86 92.64 1.65 50.79 0.46
TabDDPM+DCM 20% 0.82 90.63 0.31 94.75 2.10 44.45 0.83 98.03 0.50 94.79 1.88 91.01 2.39 50.85 0.24
@@ A.2 Experiments on More Datasets
cell21 = r=5,
cell71 = r=5,
cell121 = r=5,
hline1-2,7,12 = -,
& Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Cardio
TabDDPM 0.18 - 80.24 0.78 99.14 0.15 49.11 0.17 99.61 0.03 98.95 0.30 99.43 0.55 49.91 0.36
TabDDPM+SMOTE 0.69 79.47 0.45 96.35 2.38 47.44 1.17 96.58 0.75 94.39 1.28 85.43 1.69 50.73 0.19
TabDDPM+TCM 0.38 79.71 0.58 97.82 2.05 48.37 1.11 98.66 1.35 95.86 3.43 96.31 0.42 49.24 1.58
TabDDPM+TCMP 0.34 79.82 0.27 98.71 0.49 48.87 0.34 98.88 0.62 98.67 0.20 96.31 0.42 49.34 0.38
TabDDPM+DCM 0.24 79.41 0.46 98.92 0.62 48.84 0.14 99.19 0.36 98.72 0.32 97.38 0.65 50.93 0.40
Wilt
TabDDPM 0.35 - 99.32 0.58 98.63 0.73 50.53 0.47 98.58 1.51 98.48 0.35 98.63 1.68 52.47 0.54
TabDDPM+SMOTE 0.41 99.22 0.54 79.49 0.78 43.76 1.37 91.09 0.50 88.04 4.73 76.66 2.61 50.29 0.28
TabDDPM+TCM 0.12 99.22 0.38 97.93 1.01 48.47 1.17 97.31 1.28 95.71 2.49 96.92 1.72 48.75 2.18
TabDDPM+TCMP 0.67 99.52 0.37 98.55 0.14 49.36 0.99 97.12 0.84 96.81 0.63 96.76 3.44 45.52 1.35
TabDDPM+DCM 0.32 98.49 0.42 98.84 0.25 47.00 0.40 97.02 0.88 95.75 2.18 93.62 0.67 51.08 0.53
@@ A.3 Hyperparameter Study on Pooling Strategy
cell21 = r=4,
cell61 = r=4,
cell101 = r=4,
cell141 = r=4,
cell181 = r=4,
hline1-2,6,10,14,18 = -,
& Methods Mem. Ratio (%) Improv. MLE (%) -Precision(%) -Recall(%) Shape Score(%) Trend Score(%) C2ST(%) DCR(%)
Default
TabDDPM 0.45 - 76.79 0.69 98.15 1.45 44.41 0.70 97.58 0.95 94.46 0.68 91.85 6.04 49.12 0.94
TabDDPM+DCM-Mean 0.78 79.55 0.70 95.92 3.21 38.94 0.98 96.21 1.33 90.23 0.88 87.35 4.42 50.64 0.94
TabDDPM+DCM-Max 1.22 80.09 0.55 91.08 7.60 37.93 3.26 94.59 3.32 90.51 3.52 88.47 7.36 49.32 2.63
TabDDPM+DCM-Top-Mean 2.06 76.68 0.76 91.97 1.78 37.89 0.84 93.62 2.78 89.30 3.02 87.63 2.15 48.17 2.59
Shoppers
TabDDPM 0.31 - 92.17 0.32 93.16 1.58 52.57 1.30 97.08 0.46 92.92 3.27 86.74 0.63 51.36 0.63
TabDDPM+DCM-Mean 1.12 92.58 0.42 95.66 0.61 48.54 0.70 94.12 2.01 91.47 1.07 82.01 2.73 51.00 2.91
TabDDPM+DCM-Max 2.75 93.21 0.32 91.93 6.24 47.97 3.22 94.34 4.12 90.50 5.55 80.66 11.39 49.59 1.86
TabDDPM+DCM-Top-Mean 2.83 91.46 0.74 93.25 3.87 46.15 4.42 92.32 5.00 89.11 4.95 80.56 4.81 49.74 3.34
@@ Appendix C Experimental Details
We implement DynamicCut and DynamicCutMix along with all baseline methods using the PyTorch framework. All models are trained using the Adam optimizer.
@@ C.4 Datasets
We conduct our experiments on three widely recognized benchmark datasets sourced from the UCI Machine Learning Repository. These datasets span diverse application domains, including income prediction, credit risk assessment, and online customer behavior analysis. Table LABEL:tab:statistics provides an overview of their key characteristics.
@@
| Dataset | #Rows | #Num | #Cat | #Train | #Validation | #Test | Task |
| Adult | 48,842 | 6 | 9 | 28,943 | 3,618 | 16,281 | Classification |
| Default | 30,000 | 14 | 11 | 24,000 | 3,000 | 3,000 | Classification |
| Shoppers | 12,330 | 10 | 8 | 9,864 | 1,233 | 1,233 | Classification |
| Cardio | 70,000 | 5 | 7 | 44,800 | 11,200 | 14,000 | Classification |
| Wilt | 4,839 | 5 | 1 | 3,096 | 775 | 968 | Classification |

### v2

@@ 6.3 DynamicCut with TabCutMix (DynamicCutMix): Combined Effect and Evaluation
| Tagging | Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Default |
- | CTGAN | 0.63 | - | 68.68 0.21 | 68.95 1.70 | 16.49 0.55 | 85.04 0.93 | 77.05 2.47 | 61.89 4.89 | 49.45 0.80 |
| CTGAN | CTGAN+DCM | 0.07 | 72.37 0.39 | 71.72 0.74 | 16.64 0.20 | 85.78 0.70 | 76.85 1.62 | 57.98 3.09 | 48.48 0.24 | ||
| TabDDPM | CTGAN+DCM | 0.32 | 71.41 0.89 | 68.93 0.94 | 16.09 0.59 | 84.42 1.03 | 75.38 1.54 | 56.85 4.31 | 50.49 1.12 | ||
| - | TVAE | 0.43 | - | 72.24 0.36 | 82.97 0.37 | 20.57 0.42 | 89.37 0.54 | 83.36 0.99 | 52.63 0.37 | 51.33 0.96 | |
| TVAE | TVAE+DCM | 0.21 | 76.27 0.25 | 85.49 0.41 | 21.66 0.19 | 90.36 0.25 | 86.76 5.13 | 54.81 0.79 | 51.07 1.35 | ||
| TabDDPM | TVAE+DCM | 0.27 | 73.82 0.40 | 80.89 1.12 | 21.47 1.28 | 89.58 0.19 | 78.24 1.28 | 53.27 4.92 | 50.92 2.49 | ||
Shoppers |
- | CTGAN | 1.15 | - | 83.22 1.30 | 85.26 5.17 | 26.67 1.66 | 77.73 0.50 | 86.48 0.69 | 67.18 5.84 | 48.54 0.74 |
| CTGAN | CTGAN+DCM | 0.56 | 85.51 0.46 | 83.51 1.60 | 26.19 0.82 | 78.23 0.30 | 86.36 0.47 | 68.85 2.45 | 50.64 1.11 | ||
| TabDDPM | CTGAN+DCM | 0.43 | 83.52 0.53 | 80.21 2.75 | 25.70 1.22 | 77.49 1.24 | 86.40 0.62 | 66.03 7.67 | 50.47 2.27 | ||
| - | TVAE | 0.77 | - | 86.65 0.48 | 50.32 2.98 | 11.16 1.33 | 74.75 0.92 | 77.56 1.11 | 21.29 3.28 | 43.31 0.65 | |
| TVAE | TVAE+DCM | 1.33 | 88.32 0.29 | 51.12 2.46 | 11.98 0.58 | 75.02 0.97 | 79.02 0.12 | 22.99 2.12 | 49.98 2.55 | ||
| TabDDPM | TVAE+DCM | 0.95 | 88.43 0.17 | 50.43 3.45 | 12.08 2.01 | 74.82 0.25 | 78.37 1.18 | 22.57 3.68 | 47.26 1.81 | ||
Adult |
- | CTGAN | 0.62 | - | 88.64 0.32 | 78.15 3.66 | 26.27 0.67 | 82.43 0.90 | 82.83 0.93 | 63.64 2.74 | 49.14 0.27 |
| CTGAN | CTGAN+DCM | 0.88 | 88.45 0.29 | 79.31 1.04 | 23.31 1.42 | 82.23 0.23 | 82.80 2.27 | 63.46 3.43 | 48.13 0.33 | ||
| TabDDPM | CTGAN+DCM | 0.32 | 87.87 0.37 | 72.40 4.17 | 25.56 1.45 | 81.42 1.91 | 81.15 1.23 | 58.84 2.46 | 50.00 0.19 | ||
| - | TVAE | 0.41 | - | 88.61 0.49 | 92.34 2.19 | 29.90 1.00 | 82.71 0.45 | 79.06 0.90 | 48.67 2.70 | 48.76 0.25 | |
| TVAE | TVAE+DCM | 0.45 | 88.12 0.32 | 87.54 1.24 | 31.54 0.38 | 83.45 0.84 | 78.57 0.68 | 54.85 5.05 | 48.81 2.10 | ||
| TabDDPM | TVAE+DCM | 0.21 | 89.23 0.38 | 85.50 4.09 | 30.23 3.81 | 82.36 1.96 | 78.80 2.50 | 54.48 8.01 | 49.57 0.35 |
@@ A.1.1 Overall Performance Comparison
| Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Shoppers |
TabDDPM | 0.31 | - | 92.17 0.32 | 93.16 1.58 | 52.57 1.30 | 97.08 0.46 | 92.92 3.27 | 86.74 0.63 | 51.36 0.63 |
| TabDDPM+DCM 5% | 1.76 | 93.30 0.22 | 94.49 1.87 | 49.07 2.35 | 95.98 1.14 | 92.04 1.81 | 84.66 3.35 | 51.28 1.28 | ||
| TabDDPM+DCM 10% | 2.83 | 91.46 0.74 | 93.25 3.87 | 46.15 4.42 | 92.32 5.00 | 89.11 4.95 | 80.56 4.81 | 49.74 3.34 | ||
| TabDDPM+DCM 20% | 0.40 | 91.58 0.51 | 91.66 2.77 | 52.98 1.99 | 95.30 1.88 | 91.83 1.35 | 82.69 4.29 | 52.44 0.33 | ||
Adult |
TabDDPM | 0.18 | - | 91.09 0.07 | 93.58 1.99 | 51.52 2.29 | 98.84 0.03 | 97.78 0.07 | 94.63 1.19 | 51.56 0.34 |
| TabDDPM+DCM 5% | 1.41 | 90.91 0.26 | 95.73 2.92 | 45.56 2.63 | 98.08 0.48 | 96.14 1.50 | 90.21 3.35 | 51.06 0.56 | ||
| TabDDPM+DCM 10% | 0.57 | 90.81 0.16 | 95.23 1.14 | 45.76 1.02 | 98.27 0.32 | 96.71 0.86 | 92.64 1.65 | 50.79 0.46 | ||
| TabDDPM+DCM 20% | 0.82 | 90.63 0.31 | 94.75 2.10 | 44.45 0.83 | 98.03 0.50 | 94.79 1.88 | 91.01 2.39 | 50.85 0.24 |
@@ A.2 Experiments on More Datasets
| Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Cardio |
TabDDPM | 0.18 | - | 80.24 0.78 | 99.14 0.15 | 49.11 0.17 | 99.61 0.03 | 98.95 0.30 | 99.43 0.55 | 49.91 0.36 |
| TabDDPM+SMOTE | 0.69 | 79.47 0.45 | 96.35 2.38 | 47.44 1.17 | 96.58 0.75 | 94.39 1.28 | 85.43 1.69 | 50.73 0.19 | ||
| TabDDPM+TCM | 0.38 | 79.71 0.58 | 97.82 2.05 | 48.37 1.11 | 98.66 1.35 | 95.86 3.43 | 96.31 0.42 | 49.24 1.58 | ||
| TabDDPM+TCMP | 0.34 | 79.82 0.27 | 98.71 0.49 | 48.87 0.34 | 98.88 0.62 | 98.67 0.20 | 96.31 0.42 | 49.34 0.38 | ||
| TabDDPM+DCM | 0.24 | 79.41 0.46 | 98.92 0.62 | 48.84 0.14 | 99.19 0.36 | 98.72 0.32 | 97.38 0.65 | 50.93 0.40 | ||
Wilt |
TabDDPM | 0.35 | - | 99.32 0.58 | 98.63 0.73 | 50.53 0.47 | 98.58 1.51 | 98.48 0.35 | 98.63 1.68 | 52.47 0.54 |
| TabDDPM+SMOTE | 0.41 | 99.22 0.54 | 79.49 0.78 | 43.76 1.37 | 91.09 0.50 | 88.04 4.73 | 76.66 2.61 | 50.29 0.28 | ||
| TabDDPM+TCM | 0.12 | 99.22 0.38 | 97.93 1.01 | 48.47 1.17 | 97.31 1.28 | 95.71 2.49 | 96.92 1.72 | 48.75 2.18 | ||
| TabDDPM+TCMP | 0.67 | 99.52 0.37 | 98.55 0.14 | 49.36 0.99 | 97.12 0.84 | 96.81 0.63 | 96.76 3.44 | 45.52 1.35 | ||
| TabDDPM+DCM | 0.32 | 98.49 0.42 | 98.84 0.25 | 47.00 0.40 | 97.02 0.88 | 95.75 2.18 | 93.62 0.67 | 51.08 0.53 |
@@ A.3 Hyperparameter Study on Pooling Strategy
| Methods | Mem. Ratio (%) | Improv. | MLE (%) | -Precision(%) | -Recall(%) | Shape Score(%) | Trend Score(%) | C2ST(%) | DCR(%) | |
Default |
TabDDPM | 0.45 | - | 76.79 0.69 | 98.15 1.45 | 44.41 0.70 | 97.58 0.95 | 94.46 0.68 | 91.85 6.04 | 49.12 0.94 |
| TabDDPM+DCM-Mean | 0.78 | 79.55 0.70 | 95.92 3.21 | 38.94 0.98 | 96.21 1.33 | 90.23 0.88 | 87.35 4.42 | 50.64 0.94 | ||
| TabDDPM+DCM-Max | 1.22 | 80.09 0.55 | 91.08 7.60 | 37.93 3.26 | 94.59 3.32 | 90.51 3.52 | 88.47 7.36 | 49.32 2.63 | ||
| TabDDPM+DCM-Top-Mean | 2.06 | 76.68 0.76 | 91.97 1.78 | 37.89 0.84 | 93.62 2.78 | 89.30 3.02 | 87.63 2.15 | 48.17 2.59 | ||
Shoppers |
TabDDPM | 0.31 | - | 92.17 0.32 | 93.16 1.58 | 52.57 1.30 | 97.08 0.46 | 92.92 3.27 | 86.74 0.63 | 51.36 0.63 |
| TabDDPM+DCM-Mean | 1.12 | 92.58 0.42 | 95.66 0.61 | 48.54 0.70 | 94.12 2.01 | 91.47 1.07 | 82.01 2.73 | 51.00 2.91 | ||
| TabDDPM+DCM-Max | 2.75 | 93.21 0.32 | 91.93 6.24 | 47.97 3.22 | 94.34 4.12 | 90.50 5.55 | 80.66 11.39 | 49.59 1.86 | ||
| TabDDPM+DCM-Top-Mean | 2.83 | 91.46 0.74 | 93.25 3.87 | 46.15 4.42 | 92.32 5.00 | 89.11 4.95 | 80.56 4.81 | 49.74 3.34 |
@@ Appendix C Experimental Details
We implement DynamicCut and DynamicCutMix along with all baseline methods using the PyTorch framework. All models are trained using the Adam optimizer. Our code is available at https://anonymous.4open.science/r/DynamicCut-8255
@@ C.3 Quantifying Memorization
| Dataset | #Rows | #Num | #Cat | #Train | #Validation | #Test | Task |
| Adult | 48,842 | 6 | 9 | 28,943 | 3,618 | 16,281 | Classification |
| Default | 30,000 | 14 | 11 | 24,000 | 3,000 | 3,000 | Classification |
| Shoppers | 12,330 | 10 | 8 | 9,864 | 1,233 | 1,233 | Classification |
| Cardio | 70,000 | 5 | 7 | 44,800 | 11,200 | 14,000 | Classification |
| Wilt | 4,839 | 5 | 1 | 3,096 | 775 | 968 | Classification |
@@ C.4 Datasets
We conduct our experiments on three widely recognized benchmark datasets sourced from the UCI Machine Learning Repository. These datasets span diverse application domains, including income prediction, credit risk assessment, and online customer behavior analysis. Table 10 provides an overview of their key characteristics.